NvidiaIntely (nvidiaintely):

You wish to turn a 0.25-liter sample of 6M aqueous nitric acid (HNO3) into 3M nitric acid solution. Describe in words how you will accomplish this.

NvidiaIntely (nvidiaintely):

@Nnesha @Cas_fangirl_14 @Kevin @cuanchi

OpenStudy (mathmate):

This is chemistry. In the future, please post in the chemistry section. If there is no appropriate answer, you can post anywhere else with a \(link\) to the original post so that all discussions and solutions are in one place.

OpenStudy (cas_fangirl_14):

You would need to dilute the nitric acid with water. Add 250 mL of water. 0.25 L * 6 M = x L * 3 M 1.5 = 3x x = 0.5 L 0.5 L - 0.25 L = 0.25 L = 250 mL

OpenStudy (mathmate):

The solution to the problem is dilution by water. The amount of water to add is approximately 0.25 litres. If it is known that the original solution is exactly 6M, then use a volumetric flask to measure 250 ml of the acid and empty into a 500 ml volumetric flask. Use distilled water to rince the 250 ml flask and empty into the larger flask. Finally, add distilled water to about 1 cm below the mark, shake vigorously to make sure the mixture is well mixed, and add more distilled water to the 500 ml mark. Shake again. If original molarity is not known , use titration.

NvidiaIntely (nvidiaintely):

Thanks @mathmate!

OpenStudy (mrnood):

Incidently - in the lab always add acid to water NOT water to acid. You can cause boiling and serious acid burns if you do the oth eway!!!!

OpenStudy (mathmate):

Very true, fortunately this is nitric acid. Much heat would have been generated if it were sulphuric. 6M is more than 1:1 concentrated, and still a lot of heat would have been generated.

OpenStudy (mrnood):

This quote is fromt eh DOW chemical munufacturer's data sheet for Nitric acid: Nitric acid reacts violently with many common organic solvents and chemicals, such as ethanol, turpentine, bases, most metals, and many other materials. The reactions with organics frequently result in the formation of gas which produces additional hazards in closed systems. Contact with combustible materials such as wood chips, sawdust, cellulose and polymeric absorbents can cause a fire. Water added to nitric acid will generate heat, fumes, and splattering. If it is necessary to mix with water, nitric acid should be added slowly while stirring.